# Large Language Models in Argument Mining: A Survey

### v3

@@ 1 Introduction
Despite its clear value, manual argument annotation remains labor-intensive and inherently unscalable: crafting detailed argument maps can consume hundreds or even thousands of hours for relatively short texts. To cope with the volume of data generated online, research in Argument Mining (AM) has accelerated significantly since around 2013, marked by the launch of dedicated workshops at ACL and themed tutorials at top-tier conferences including IJCAI, ACL, and ESSLLI (Cabrio and Villata, 2018). Foundational surveys—most notably those by Lawrence and Reed (2019), which comprehensively cataloged AM techniques, datasets, and challenges (Lawrence and Reed, 2019); Peldszus and Stede (2013), who systematically bridged classical argument-diagramming methods with machine-readable annotation pipelines (Peldszus and Stede, 2013); and Lytos et al. (2019), who documented AM’s adaptation to less formal, user-generated text (Lytos et al., 2019)—have provided essential context for early developments.
@@
However, despite this breadth, most existing reviews remain anchored in the pre-LLM era. They emphasize supervised learning and feature engineering, formal annotation schemas, or segmented task pipelines, and so far have only cursorily addressed the revolutionary role of pre-trained language models. In response, a new wave of empirical literature has emerged—exemplified by work on LLM-driven argument summarization, argument extraction, and relationship classification (Chen et al., 2024; Altemeyer et al., 2025; Guida et al., 2025). Yet, there is still no review that systematically integrates these advances, examines how LLMs reshape the boundaries and synergies of AM subtasks, examining how advances in one area reshape assumptions in another, and reconnects these technical developments to foundational argumentation theory and annotation practice. We treat argument mining as an interlocking system whose components—tasks, data, metrics, and models—evolve together under the pressure of foundation-model capabilities.
@@ 3.3 Relation Identification
If claim detection marks the nodes of an argumentative graph, relation identification supplies its edges. Early work treated the task as sentence-pair classification—“support”, “attack”, or “unrelated”—in tightly controlled genres such as German news commentary (Eckle-Kohler et al., 2015). Genre expansion quickly revealed edge sparsity and annotation drift: the AMPERSAND corpus of Reddit threads, for example, required additional labels like rebuttal and mitigation to capture conversational nuance (Chakrabarty et al., 2020). Methodologically, the field has moved from feature-rich SVMs to graph-neural encoders that propagate information across discourse trees (DBLP:conf/acl/ZhangNL23). Yet even state-of-the-art models stumble when relations span paragraphs, rely on implicit premises, or cross language boundaries. .
@@ 5.2 Modelling Advances: From Feature Engineering to LLM-based Judges
The 2016–2021 period marks an unequivocal shift from sparse inputs to representation learning. Bidirectional LSTMs outperformed bag-of-words on convincingness by capturing sequential cues (Habernal and Gurevych, 2016); CNN–BiLSTM hybrids added early cross-topic generalization on comment threads (Huber et al., 2019). Several groups then moved from classification to ranking losses: margin-based pairwise objectives (Potash et al., 2019) and listwise NDCG targets (Gu et al., 2018) yielded more stable quality orderings. Transformer fine-tuning largely removed feature design: BERT and RoBERTa lifts of 5–10 F1 or $\rho$ on IBM-Rank and GAQ tasks are now well documented (Gretz et al., 2020; Lauscher et al., 2020). Downstream improvements came from metric-learning variants—e.g. Siamese-SBERT with contrastive loss for out-of-domain convincingness (DBLP:conf/argmining/ZhangR23)—and from multi-task heads that jointly predict stance or component type, boosting macro-F1 on low-resource sets (Favreau et al., 2022). Efficient pairwise crowdsourcing accelerated data creation and ranking-loss training (Gienapp et al., 2020), while contextual embeddings enabled intrinsic, reference-free scoring of short arguments (Wachsmuth and Werner, 2020). Together, these advances consolidated neural encoders as the default for AQA.
@@ 6.4 Synthetic Supervision and Data Augmentation
The downside is distributional drift. Synthetic arguments tend to be lexically fluent, hedge-free and uniformly polite, properties that can over-regularise fine-tuned models. Provenance cards documenting prompt templates, sampling temperatures, rejection rates and post-editing guidelines are therefore being proposed as a best practice (DBLP:conf/facc/Paullada2021). Legal uncertainty over copyright of machine-generated text remains unresolved.

### v4

@@ 1 Introduction
Despite its clear value, manual argument annotation remains labor-intensive and inherently unscalable: crafting detailed argument maps can consume hundreds or even thousands of hours for relatively short texts. To cope with the volume of data generated online, research in Argument Mining (AM) has accelerated significantly since around 2013, marked by the launch of dedicated workshops at ACL and themed tutorials at top-tier conferences including IJCAI, ACL, and ESSLLI (Cabrio and Villata, 2018). Foundational surveys—most notably those by Lawrence and Reed (2019), which comprehensively cataloged AM techniques, datasets, and challenges (Lawrence and Reed, 2019); Patel Peldszus and Stede (2013), who systematically bridged classical argument-diagramming methods with machine-readable annotation pipelines (Peldszus and Stede, 2013); and Lytos et al. (2019), who documented AM’s adaptation to less formal, user-generated text (Lytos et al., 2019)—have provided essential context for early developments.
Figure 1: Key large language models and techniques influencing argument mining.
@@
However, despite this breadth, most existing reviews remain anchored in the pre-LLM era. They emphasize supervised learning and feature engineering, formal annotation schemas, or segmented task pipelines, and so far have only cursorily addressed the revolutionary role of pre-trained language models. In response, a new wave of empirical literature has emerged, exemplified by work on LLM-driven argument summarization, argument extraction, and relationship classification (Chen et al., 2024; Altemeyer et al., 2025; Guida et al., 2025). Yet, there is still no review that systematically integrates these advances (shown in Figure 1), examines how LLMs reshape the boundaries and synergies of AM subtasks, examining how advances in one area reshape assumptions in another, and reconnects these technical developments to foundational argumentation theory and annotation practice. We treat argument mining as an interlocking system whose components—tasks, data, metrics, and models evolve together under the pressure of foundation-model capabilities.
@@ 3.3 Relation Identification
If claim detection marks the nodes of an argumentative graph, relation identification supplies its edges. Early work treated the task as sentence-pair classification—“support”, “attack”, or “unrelated”—in tightly controlled genres such as German news commentary (Eckle-Kohler et al., 2015). Genre expansion quickly revealed edge sparsity and annotation drift: the AMPERSAND corpus of Reddit threads, for example, required additional labels like rebuttal and mitigation to capture conversational nuance (Chakrabarty et al., 2020). Yet even state-of-the-art models stumble when relations span paragraphs, rely on implicit premises, or cross language boundaries. .
@@ 5.2 Modelling Advances: From Feature Engineering to LLM-based Judges
The 2016–2021 period marks an unequivocal shift from sparse inputs to representation learning. Bidirectional LSTMs outperformed bag-of-words on convincingness by capturing sequential cues (Habernal and Gurevych, 2016); CNN–BiLSTM hybrids added early cross-topic generalization on comment threads (Huber et al., 2019). Several groups then moved from classification to ranking losses: margin-based pairwise objectives (Potash et al., 2019) and listwise NDCG targets (Gu et al., 2018) yielded more stable quality orderings. Transformer fine-tuning largely removed feature design: BERT and RoBERTa lifts of 5–10 F1 or $\rho$ on IBM-Rank and GAQ tasks are now well documented (Gretz et al., 2020; Lauscher et al., 2020). Downstream improvements came from metric-learning variants—e.g. Siamese-SBERT with contrastive loss for out-of-domain convincingness and from multi-task heads that jointly predict stance or component type, boosting macro-F1 on low-resource sets (Favreau et al., 2022). Efficient pairwise crowdsourcing accelerated data creation and ranking-loss training (Gienapp et al., 2020), while contextual embeddings enabled intrinsic, reference-free scoring of short arguments (Wachsmuth and Werner, 2020). Together, these advances consolidated neural encoders as the default for AQA.
@@ 6.4 Synthetic Supervision and Data Augmentation
The downside is distributional drift. Synthetic arguments tend to be lexically fluent, hedge-free and uniformly polite, properties that can over-regularise fine-tuned models. Provenance cards documenting prompt templates, sampling temperatures, rejection rates and post-editing guidelines are therefore being proposed as a best practice . Legal uncertainty over copyright of machine-generated text remains unresolved.